# Learning dynamical systems from data: Gradient-based dictionary optimization

### v1

@@ 1 Introduction
The accuracy of the learned transfer operators or governing equations depends strongly on the dictionary. Poorly chosen basis functions can result in ill-conditioned matrices, spectral pollution, and incorrect predictions. Selecting suitable basis functions is an open problem. In practice, often dictionaries comprising, for instance, monomials, trigonometric functions, or radial-basis functions are used. Our goal is to develop a more flexible framework for data-driven modeling approaches that not only learns the dynamics but also, at the same time, the basis functions. While neural network-based dictionary learning methods have been successfully used for approximating transfer operators and to identify governing equations [15, 16, 17, 18], we lose the interpretability of the dictionary and cannot immediately identify the governing equations. We propose a dictionary learning approach that is based on well-known gradient-based optimization techniques such as stochastic gradient descent (SGD), Nesterov’s method, or Adam, see [19, 20, 21], and allows us to optimize the parameters of basis functions (e.g., the centers and bandwidths of Gaussians). The key advantage of our method is that we obtain an interpretable representation. It can be regarded as a compromise between the flexibility of fully optimizable basis functions generated by a neural network and the interpretability and expressivity of pre-selected dictionaries. Alternating optimization algorithms to learn the Koopman operator have also been proposed in [22, 16] using the reconstruction error as a loss function. However, we present a general framework for alternating optimization to learn the dynamics and optimal and interpretable dictionaries using different loss functions. The main contributions of this work are:
@@
The remainder of the paper is structured as follows: We first introduce the required concepts including transfer operators, EDMD, and SINDy as well as gradient-based optimization algorithms in Section 2. In Section 3, we discuss the proposed optimization framework for learning dynamical systems from data. We present the numerical results for various deterministic and stochastic dynamical systems in Section 4. Open problems and a discussion of future work can be found in Section 5.
@@ 2.2 Extended dynamic mode decomposition
EDMD aims to find the best approximation of the Koopman operator $\mathcal{K}^{\tau}$ projected onto the space spanned by the selected basis functions. In what follows, let $\|\cdot\|_{F}$ denote the Frobenius norm. In order to obtain an approximation of the Koopman operator, we first map the training data to the feature space by defining
@@
$K^{\top}=\Psi_{y}\Psi_{x}^{+}=(\Psi_{y}\Psi_{x}^{\top})(\Psi_{x}\Psi_{x}^{\top})^{+}=C_{yx}\hskip 1.00006ptC_{xx}^{+},$
where + denotes the pseudoinverse and $C_{xx}$ and $C_{xy}$ are the (uncentered) covariance and cross-covariance matrices, respectively. The matrix $K\in\mathbb{R}^{n\times n}$ is the matrix representation of the projected Koopman operator $\mathcal{K}^{\tau}$. We can then compute the eigenvalues $\lambda_{i}$ and the eigenvectors $v_{i}$ of the matrix $K$ to obtain eigenfunctions
@@ 2.3 Sparse identification of nonlinear dynamics
Given training data of the form $\{(x_{i},\dot{x}_{i})\}_{i=1}^{m}$, which can be, for example, extracted from one trajectory, we construct the data matrix
@@
The matrix $\dot{X}$ contains the time-derivatives at the training data points, which can, for example, be approximated numerically using finite differences. In order to find an approximation of the governing equations, we minimize the loss function
@@ 2.4 SINDy for PDEs
where the subscripts denotes the partial derivatives with respect to $t$ and $x$. Suppose that we have data from a solution of the PDE on an $n\times m$ grid, i.e., for $m$ spatial and $n$ time points. We start by constructing a discretized version of the right-hand side of the above equation and express $u(x,t)$ and its derivatives using the matrix $\Theta(U)$. Each column of the matrix $\Theta(U)$ represents a candidate term that can be present in the right-hand side of the PDE. That is, if we have $d$ candidate terms, then there will be $d$ columns. Each column represents the value of a particular candidate function on all $mn$ space-time points. We can write the discretized version of the PDE as
$U_{t}=\Theta(U)\xi,$
where $\Theta(U)\in\mathbb{R}^{mn\times d}$. Each nonzero entry in $\xi$ corresponds to a term in the PDE selected from the dictionary of candidate terms. To obtain a sparse vector, the algorithm minimizes the following loss function
@@ 2.5 Gradient descent algorithms
$x^{*}=\operatorname*{argmin}\limits_{x\in\mathbb{R}^{n}}Q(x).$
There exist many different gradient descent algorithms to find the minimizer $x^{*}$ of the above unconstrained optimization problem. We will briefly review some of these methods, but refer the reader to [20, 21, 25, 26] for more details.
@@ 2.5.1 Gradient descent
Given a loss function (also sometimes called a cost function or objective function) $Q$, gradient descent (GD) aims to find a minimizer by following the gradient “downhill”. Given an initial guess $x_{0}$, GD is defined by
@@ 2.5.2 Stochastic gradient descent
We then approximate the actual gradient $\nabla_{x}Q(x)$ by the gradient for a single data point $\nabla_{x}Q_{i}(x)$ and compute
@@
For optimization problems involving a large number of data points, SGD reduces the computational costs considerably. However, the convergence can be very slow due to the approximation of the true gradient. A compromise between computational efficiency and accuracy is to compute the gradient with respect to a random subset or batch of data points. This is known as mini-batch SGD.
@@ 2.5.4 Adam
The last optimization method that we will consider is Adam [19], which is an extension of SGD. It is among the most popular methods for training neural networks as it is computationally efficient for large-scale optimization problems, has reasonable memory requirements, and is suitable for noisy gradients as well. The algorithm is based on the adaptive estimation of first- and second-order moments and involves the calculation of moving averages of the gradients while controlling the decay of these averages with the parameters $\beta_{1}$ and $\beta_{2}$. The values of the hyperparameters are typically set to the default values $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1\times 10^{-8}$. The main steps for the algorithm are:
@@
  • Update rule of Adam: $x_{t+1}=x_{t}-h\cdot\hat{m}_{t+1}/(\sqrt{\hat{v}_{t+1}}+\epsilon)$.
@@ 3.1 Parametric EDMD
using gradient descent algorithms to find $K$. However, the reconstruction error might not be a good choice for optimizing the parameters of the basis functions. Suppose, for example, we choose the basis functions to be Gaussians with variable centers and bandwidths. In that case, the bandwidths of the Gaussian functions might tend to infinity so that $\Psi_{x}(w)$ and $\Psi_{y}(w)$ become almost constant and choosing $K=I$ minimizes the loss function. To avoid this problem, we utilize the variational approach for Markov processes (VAMP) score. The VAMP-2 score is defined by
@@
For the loss function $\mathcal{F}(K,w)=\big{\|}\Psi_{y}(w)-K^{\top}\Psi_{x}(w)\big{\|}_{F}^{2}$, we write $\mathcal{F}(K,w)=\sum_{i=1}^{m}\mathcal{F}_{i}(K,w)$, where
$\mathcal{F}_{i}(K,w)=\big{\|}\psi(y_{i},w)-K^{\top}\psi(x_{i},w)\big{\|}_{2}^{2}$
and $m$ is the number of data points.
@@
This follows immediately from $\|A\|_{F}^{2}=\sum_{i}\|a_{i}\|_{2}^{2}$, where $a_{i}$ denotes the $i$th column of $A$. ∎
@@
with $A=\Psi_{x}^{\top}(w)\in\mathbb{R}^{m\times n}$ and $b=[\Psi_{y}^{\top}(w)]_{i}\in\mathbb{R}^{m}$. The application of gradient-descent techniques to regression problems of this form has been studied, for example, in [21]. It has been shown that a good choice of step size is the inverse of the largest eigenvalue of the matrix $A^{\top}A$. We thus choose the step size $h=\frac{1}{\lambda_{max}}$, where $\lambda_{max}$ is the largest eigenvalue of $\Psi_{x}(w)\Psi_{x}^{\top}(w)$.
@@ 3.2 Parametric SINDy for a system of ODEs
The cost function $\mathcal{F}(\Xi,w)=\big{\|}\dot{X}-\Xi^{\top}\Psi_{x}(w)\big{\|}_{F}^{2}$ can be written as $\mathcal{F}(\Xi,w)=\sum_{i=1}^{m}\mathcal{F}_{i}(\Xi,w)$, where $\mathcal{F}_{i}(\Xi,w)=\big{\|}\dot{x_{i}}-\Xi^{\top}\psi(x_{i},w)\big{\|}_{2}^{2}$ and $m$ is the number of data points.
@@ 3.3 Parametric SINDy for PDEs
where $N(\cdot)$ is a nonlinear function of $u(x,t)$ and its partial derivatives with respect to $x$ and $t$ and $w=[w_{1},\ldots,w_{p}]^{\top}\in\mathbb{R}^{p}$ is the vector of parameters of the PDE. We create the matrix $\Theta(U(w))\in\mathbb{R}^{mn\times d}$ of the basis functions as discussed in Section 2. We then minimize the loss function
$\mathcal{F}(\xi,w)=\big{\|}U_{t}-\Theta(U(w))\xi\big{\|}_{2}^{2}+\epsilon\hskip 1.00006pt\kappa(\Theta(U(w)))$ (6)
@@ 3.4 Proposed algorithm
We now illustrate the proposed optimization framework. Let $\mathcal{L}_{1}(A,w)$ and $\mathcal{L}_{2}(A,w)$ be two arbitrary loss functions. We define an alternating Adam algorithm as shown in Algorithm 1. This allows us to use the implementation for both parametric EDMD and SINDy. Other gradient-based optimization methods can be implemented in a similar way.
@@
  • $m^{a}_{0}=0$ (first moment for $A$) & $v^{k}_{0}=0$ (second moment for $A$)
@@ 4.1.1 Protein folding
Analyzing the structure of proteins, especially transitions between folded and unfolded states, helps us understand biological processes and develop new drugs or treatments [31]. We consider the protein Chignolin (CLN025) consisting of $10$ residues and $166$ atoms. The simulation data was generated by D.E. Shaw Research, see [32] for more details. The trajectory is $1.06\cdot 10^{8}$ ps long. We subsample the trajectory to create the data $\{x_{i},y_{i}\}_{i=1}^{m}$ using the lag time $\tau=9911.5$ ps. We then transform the trajectory data using contact maps, which are symmetric matrices containing the distances between different pairs of residues, resulting in $\Psi_{x},\Psi_{y}\in\mathbb{R}^{28\times 10,694}$. Since the basis functions are in this case the contact distances, there are no parameters to optimize.
To approximate the Koopman operator, we apply standard EDMD and the proposed optimization framework. The relative approximation error for different gradient descent-based algorithms compared to the exact EDMD solution are shown in Figure LABEL:subfig:error_chig. Eigenvalues of the approximated Koopman operator are given in Figure LABEL:subfig:chig_eigv. There is a spectral gap between the second and third eigenvalue, which indicates that there exist two metastable states. Furthermore, the values of the two dominant eigenfunctions, shown in Figure LABEL:subfig:chig_eigf_oa, can be clustered into two sets representing the folded and the unfolded states of the molecule. Examples of folded and unfolded states are shown in Figures LABEL:subfig:chig_folded_states and LABEL:subfig:chig_unfolded_states, respectively. The frequency of contacts between different residues of the molecule for the set of folded and unfolded states is presented in Figures LABEL:subfig:chig_folded_contact and LABEL:subfig:chig_unfolded_contact, respectively. We can see that Chignolin is mostly in its folded (native) state. The example shows that the gradient descent algorithms converge to the EDMD approximation and that Adam eventually outperforms the other gradient descent approaches.
@@
Figure 2: (a) & (b) Folded and unfolded state of the Chignolin protein. (c) & (d) Frequency of contacts between different residue pairs over all the identified folded and unfolded states, respectively.
@@ 4.1.2 Ornstein–Uhlenbeck process
$\lambda_{i}=e^{-\alpha(i-1)},\quad\varphi_{i}(x)=\frac{1}{\sqrt{(i-1)!}}H_{i-1}\big{(}\sqrt{\alpha\beta}x\big{)},\quad i=1,2,3,\dots,$
@@
i.e., one-dimensional Gaussian functions with variable centers $c_{i}$ and bandwidths $\sigma_{i}$. That is, $w_{i}=[c_{i},\sigma_{i}]^{\top}$ and $w\in\mathbb{R}^{2\hskip 0.70004ptn}$ contains the centers and bandwidths of all basis functions. We select the $n=14$ basis functions shown in Figure LABEL:subfig:initial_basis_ou to approximate the Koopman operator. Note that most of the centers of the Gaussian functions are initially in the left half of the domain. As a result, the approximation of the eigenfunctions will be poor in the right half of the domain. Applying the alternating optimization algorithm allows us to find more suitable basis functions. The optimized basis functions are shown in Figure LABEL:subfig:final_basis_ou. The centers are now evenly distributed and lead to a better approximation of the Koopman operator. The convergence of the reconstruction error (3) is presented in Figure LABEL:subfig:errorK_ou and the convergence of the VAMP-2 score in Figure LABEL:subfig:vamp_ou. The eigenvalues and eigenfunctions of the approximated Koopman operator are presented in Figures LABEL:subfig:ou_eigv and LABEL:subfig:ou_eigf, respectively. The numerically computed eigenvalues and eigenfunctions match the theoretical results.
Figure 3: Koopman operator approximation for the Ornstein–Uhlenbeck process using parametric Gaussian functions. (a) Initial basis functions. (b) Optimized basis functions. (c) Convergence of the reconstruction error. (d) Convergence of the VAMP-2 score for optimizing the parameters of the basis functions. (e) Three dominant eigenvalues of the Koopman operator. (f) Corresponding eigenfunctions of the Koopman operator, where the light-colored lines represent the true eigenfunctions.

### v2

@@ 1 Introduction
The accuracy of the learned transfer operators or governing equations depends strongly on the dictionary. Poorly chosen basis functions can result in ill-conditioned matrices, spectral pollution, and incorrect predictions. Selecting suitable basis functions is an open problem. In practice, often dictionaries comprising, for instance, monomials, trigonometric functions, or radial-basis functions are used. Our goal is to develop a more flexible framework for data-driven modeling approaches that not only learns the dynamics but also, at the same time, the basis functions. Although neural network-based dictionary learning methods have been successfully used for approximating transfer operators in combination with EDMD and to identify governing equations, see, e.g., [15, 16, 17, 18, 19], where the output layer of the network represents the basis functions, we lose the interpretability of the dictionary and cannot immediately identify the governing equations. We propose a dictionary learning approach that is based on well-known gradient-based optimization techniques such as stochastic gradient descent (SGD), Nesterov’s method, or Adam, see [20, 21, 22], and allows us to optimize the parameters of basis functions (e.g., the centers and bandwidths of Gaussians). The key advantage of our method is that it is a generalized framework that encapsulates different data-driven algorithms for learning dynamical systems and results in an interpretable representation. It can be regarded as a compromise between the flexibility of fully optimizable basis functions generated by a neural network and the interpretability and expressivity of pre-selected dictionaries. Alternating optimization algorithms to learn the Koopman operator have also been proposed in [23, 19] using the reconstruction error as a loss function. However, we present a general framework for alternating optimization to learn the dynamics and optimal and interpretable dictionaries using different loss functions. The main contributions of this work are:
@@
The remainder of the paper is structured as follows: We first introduce the required concepts including transfer operators, EDMD, and SINDy as well as gradient-based optimization algorithms in Section 2. In Section 3, we discuss the proposed optimization framework for learning dynamical systems from data. We present numerical results for various deterministic and stochastic dynamical systems in Section 4. Open problems and a discussion of future work can be found in Section 5.
@@ 2.2 Extended dynamic mode decomposition
EDMD aims to find the best approximation of the Koopman operator $\mathcal{K}^{\tau}$ projected onto the space spanned by the selected basis functions. In what follows, let $\lVert\cdot\rVert_{F}$ denote the Frobenius norm. In order to obtain an approximation of the Koopman operator, we first map the training data to the feature space by defining
@@
$K^{\top}=\Psi_{y}\Psi_{x}^{+}=\big{(}\Psi_{y}\Psi_{x}^{\top}\big{)}\big{(}\Psi_{x}\Psi_{x}^{\top}\big{)}^{+}=C_{yx}\hskip 1.00006ptC_{xx}^{+},$
where + denotes the pseudoinverse and $C_{xx}$ and $C_{xy}=C_{yx}^{\top}$ are the (uncentered) covariance and cross-covariance matrices, respectively. The matrix $K\in\mathbb{R}^{n\times n}$ is the matrix representation of the projected Koopman operator $\mathcal{K}^{\tau}$. We can then compute the eigenvalues $\lambda_{i}$ and the eigenvectors $v_{i}$ of the matrix $K$ to obtain eigenfunctions
@@ 2.3 Sparse identification of nonlinear dynamics
Given training data of the form $\{(x_{i},\dot{x}_{i})\}_{i=1}^{m}$, which can, for example, be extracted from one long trajectory, we construct the data matrix
@@
The matrix $\dot{X}$ contains the time-derivatives at the training data points, which can also be approximated numerically using finite differences. In order to find an approximation of the governing equations, we minimize the loss function
@@ 2.4 SINDy for PDEs
where the subscripts denote the partial derivatives with respect to $t$ and $x$. Suppose that we have data from a solution of the PDE on an $n\times m$ grid, i.e., for $m$ spatial and $n$ time points. We start by constructing a discretized version of the right-hand side of the above equation and express $u(x,t)$ and its derivatives using the matrix $\Theta(U)$. Each column of the matrix $\Theta(U)$ represents a candidate term that can be present in the right-hand side of the PDE. That is, if we have $d$ candidate terms, then there will be $d$ columns. Each column represents the value of a particular candidate function on all $mn$ space-time points. We can write the discretized version of the PDE as
$U_{t}=\Theta(U)\hskip 1.00006pt\xi,$
where $\Theta(U)\in\mathbb{R}^{mn\times d}$. Each nonzero entry in $\xi$ corresponds to a term in the PDE selected from the dictionary of candidate terms. To obtain a sparse vector, the algorithm minimizes the loss function
@@ 2.5 Gradient descent algorithms
$x^{*}=\operatorname*{argmin}\limits_{x\in\mathbb{R}^{n}}Q(x),$
where $Q$ is called loss function, cost function, or objective function. There exist many different gradient descent algorithms to find a minimizer $x^{*}$ of the above unconstrained optimization problem. We will briefly review some of these methods, but refer the reader to [21, 22, 26, 27] for more details.
@@ 2.5.1 Gradient descent
Gradient descent (GD) aims to find a minimizer by following the gradient “downhill”. Given an initial guess $x_{0}$, GD is defined by
@@ 2.5.2 Stochastic gradient descent
We then approximate the actual gradient $\nabla_{x}Q(x)$ by the gradient $\nabla_{x}Q_{i}(x)$ w.r.t. a single data point and compute
@@
For optimization problems involving a large number of data points, SGD reduces the computational costs considerably. However, the convergence can be very slow due to the approximation of the true gradient. A compromise between computational efficiency and accuracy is to compute the gradient with respect to a random subset or so-called batch of data points. This is known as mini-batch SGD.
@@ 2.5.4 Adam
The last optimization method that we will consider is Adam [20], which is an extension of SGD. It is among the most popular methods for training neural networks as it is computationally efficient for large-scale optimization problems, has reasonable memory requirements, and is suitable for noisy gradients as well. The algorithm is based on the adaptive estimation of first- and second-order moments and involves the calculation of moving averages of the gradients while controlling the decay of these averages with the parameters $\beta_{1}$ and $\beta_{2}$. For a theoretical analysis of the convergence of Adam, see, e.g., [29, 30]. The values of the hyperparameters are typically set to the default values $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1\times 10^{-8}$. The main steps for the algorithm are:
@@
  • Compute next iterate: $x_{t+1}=x_{t}-h\cdot\hat{m}_{t+1}/(\sqrt{\hat{v}_{t+1}}+\epsilon)$.
@@ 3.1 Parametric EDMD
using gradient descent algorithms to find $K$. However, the reconstruction error might not be a good choice for optimizing the parameters of the basis functions. Suppose, for example, we choose the basis functions to be Gaussians with variable centers and bandwidths. In that case, the bandwidths of the Gaussian functions might tend to infinity so that $\Psi_{x}(w)$ and $\Psi_{y}(w)$ become almost constant and choosing $K=I$ minimizes the loss function. To avoid this problem, we utilize the variational approach for Markov processes (VAMP) score [15]. The VAMP-2 score is defined by
@@
For the loss function $\mathcal{F}(K,w)=\big{\|}\Psi_{y}(w)-K^{\top}\Psi_{x}(w)\big{\|}_{F}^{2}$, we write
$\mathcal{F}(K,w)=\sum_{i=1}^{m}\mathcal{F}_{i}(K,w),\quad\text{with}\quad\mathcal{F}_{i}(K,w)=\big{\|}\psi(y_{i},w)-K^{\top}\psi(x_{i},w)\big{\|}_{2}^{2},$
where $m$ is the number of data points.
@@
This follows immediately from the property $\|A\|_{F}^{2}=\sum_{i}\|a_{i}\|_{2}^{2}$, where $a_{i}$ denotes the $i$th column of $A$. ∎
@@
with $A=\Psi_{x}^{\top}(w)\in\mathbb{R}^{m\times n}$ and $b=[\Psi_{y}^{\top}(w)]_{i}\in\mathbb{R}^{m}$. The application of gradient-descent techniques to regression problems of this form has been studied, for example, in [22]. It has been shown that a good choice for the step size is the inverse of the largest eigenvalue of the matrix $A^{\top}A$. We thus choose the step size $h=\frac{1}{\lambda_{max}}$, where $\lambda_{max}$ is the largest eigenvalue of $\Psi_{x}(w)\Psi_{x}^{\top}(w)$.
@@ 3.2 Parametric SINDy for a system of ODEs
The cost function $\mathcal{F}(\Xi,w)=\big{\|}\dot{X}-\Xi^{\top}\Psi_{x}(w)\big{\|}_{F}^{2}$ can be written as
$\mathcal{F}(\Xi,w)=\sum_{i=1}^{m}\mathcal{F}_{i}(\Xi,w),\quad\text{with}\quad\mathcal{F}_{i}(\Xi,w)=\big{\|}\dot{x_{i}}-\Xi^{\top}\psi(x_{i},w)\big{\|}_{2}^{2},$
where $m$ is the number of data points.
@@ 3.3 Parametric SINDy for PDEs
where $N(\cdot)$ is a nonlinear function of $u(x,t)$ and its partial derivatives with respect to $x$ and $t$ and $w$ is the vector of parameters of the PDE. We create the matrix $\Theta(U(w))\in\mathbb{R}^{mn\times d}$ of the basis functions as discussed in Section 2. We then minimize the loss function
$\mathcal{F}(\xi,w)=\big{\|}U_{t}-\Theta(U(w))\hskip 1.00006pt\xi\big{\|}_{2}^{2}$ (6)
@@
Proposition 3.6.
For fixed $w$, the derivative of the loss function (6) with respect to $\xi$ is given by
$\frac{\partial}{\partial\xi}\mathcal{F}(\xi,w)=2\hskip 1.00006pt\Theta(U(w))^{\top}\Theta(U(w))\hskip 1.00006pt\xi-2\hskip 1.00006pt\Theta(U(w))^{\top}U_{t}.$
Proof.
This follows from the proof of Proposition 3.1. ∎
For computing the gradient of (6) with respect to $w$, we again use JAX.
Remark 3.7.
In order to avoid overfitting, it would also be possible to add regularization terms to the various loss functions. For example, SINDy is based on the assumption that there are only a few terms governing the dynamics. We can hence include an $L^{1}$ regularization of the matrix $\Xi$ to obtain a sparse solution. For the PDE identification approach, a regularization based on the condition number of the matrix $\Theta(U(w))$ might be beneficial as shown in [14].
@@ 3.4 Proposed algorithm
We now illustrate the proposed optimization framework. Let $\mathcal{L}_{1}(A,w)$ and $\mathcal{L}_{2}(A,w)$ be two arbitrary loss functions. We define an alternating Adam algorithm, which can be used for both parametric EDMD and SINDy. Other gradient-based optimization methods can be implemented in a similar way.
@@
  • $m^{a}_{0}=0$ (first moment for $A$) & $v^{a}_{0}=0$ (second moment for $A$)
@@ 4.1.1 Protein folding
Analyzing the structure of proteins, especially transitions between folded and unfolded states, helps us understand biological processes and develop new drugs or treatments [34]. We consider the protein Chignolin (CLN025) consisting of $10$ residues and $166$ atoms. The simulation data was generated by D.E. Shaw Research, see [35] for more details. The trajectory is $1.06\cdot 10^{8}$ ps long. We subsample the trajectory to create the training data $\{x_{i},y_{i}\}_{i=1}^{m}$ using the lag time $\tau=9911.5$ ps. We then transform the trajectory data using contact maps, which measure the distances between different pairs of residues, resulting in $\Psi_{x},\Psi_{y}\in\mathbb{R}^{28\times 10,694}$. Since the basis functions are in this case the contact distances, there are no parameters to optimize.
To approximate the Koopman operator, we apply standard EDMD and the proposed optimization framework. The relative approximation error for different gradient descent-based algorithms compared to the exact EDMD solution are shown in Figure LABEL:subfig:error_chig. Eigenvalues of the approximated Koopman operator are given in Figure LABEL:subfig:chig_eigv. There is a spectral gap between the second and third eigenvalue, which indicates that there exist two metastable states. Furthermore, the values of the two dominant eigenfunctions, shown in Figure LABEL:subfig:chig_eigf_oa, can be clustered into two sets representing the folded and unfolded states of the molecule. Examples of folded and unfolded states are shown in Figures LABEL:subfig:chig_folded_states and LABEL:subfig:chig_unfolded_states, respectively. The frequency of contacts between different residues of the molecule for the set of folded and unfolded states is presented in Figures LABEL:subfig:chig_folded_contact and LABEL:subfig:chig_unfolded_contact, respectively. We can see that Chignolin is mostly in its folded (native) state. The example shows that the gradient descent algorithms converge to the EDMD approximation and that Adam eventually outperforms the other gradient descent approaches.
@@
Figure 2: (a) & (b) Folded and unfolded states of the Chignolin protein. (c) & (d) Frequency of contacts between different residue pairs over all the identified folded and unfolded states, respectively.
@@ 4.1.2 Ornstein–Uhlenbeck process
$\lambda_{i}=e^{-\alpha(i-1)\tau},\quad\varphi_{i}(x)=\frac{1}{\sqrt{(i-1)!}}H_{i-1}\big{(}\sqrt{\alpha\beta}x\big{)},\quad i=1,2,3,\dots,$
@@
i.e., Gaussian functions with variable centers $c_{i}$ and bandwidths $\sigma_{i}$. That is, $w_{i}=[c_{i},\sigma_{i}]^{\top}$ and $w\in\mathbb{R}^{2\hskip 0.70004ptn}$ contains the centers and bandwidths of all basis functions. We select the $n=14$ basis functions shown in Figure LABEL:subfig:initial_basis_ou to approximate the Koopman operator. Note that most of the centers of the Gaussian functions are initially in the left half of the domain. As a result, the approximation of the eigenfunctions will be poor in the right half. Applying the alternating optimization algorithm allows us to find more suitable basis functions. The optimized basis functions are shown in Figure LABEL:subfig:final_basis_ou. The centers are now evenly distributed and lead to a better approximation of the Koopman operator. The convergence of the reconstruction error (3) is presented in Figure LABEL:subfig:errorK_ou and the convergence of the VAMP-2 score in Figure LABEL:subfig:vamp_ou. The eigenvalues and eigenfunctions of the approximated Koopman operator are presented in Figures LABEL:subfig:ou_eigv and LABEL:subfig:ou_eigf, respectively. The numerically computed eigenvalues and eigenfunctions match the theoretical results.
Figure 3: Koopman operator approximation for the Ornstein–Uhlenbeck process using parametric Gaussian functions. (a) Initial basis functions. (b) Optimized basis functions. (c) Convergence of the reconstruction error. (d) Convergence of the VAMP-2 score for optimizing the parameters of the basis functions. (e) Three dominant eigenvalues of the Koopman operator. (f) Corresponding eigenfunctions of the Koopman operator, where the solid lines represent the true eigenfunctions.